Many businesses rent a limousine to chauffeur important clients to/from airport, hotel, or office. A random sample of the cost (in dollars) of renting a limousine for an entire day in Cincinnati was obtained. The sample size was 51, the sample mean was 410.25 and the sample standard deviation was 35.07. Let μ be the population mean of the cost of renting a limousine for businesses for an entire day in Cincinnati.

  1. Obtain a point estimate for μ and its standard error.

  2. Construct a 95% two-sided confidence interval for μ.

  3. Construct a 99% one-sided lower confidence interval for μ.

  4. Test the hypothesis μ = 400 against μ \small \neq 400 using a significance level of α = 0.05. Provide all the details including test statistic, rejection region and the decision of whether to reject H0.

  5. Test the hypothesis μ ≤ 400 against μ > 400 using a significance level of α = 0.01. Provide all the details including test statistic, rejection region and the decision of whether to reject H0.
Answer #1

Solution: Sample size n 51 Sample mean - 410.25 Sample Standard Deviation S 35.07 Degrees of freedom-df = n-1 = 51-1 = 50 Poi